Three-sevenths of a number is 21.
Find the number.
the number is 9
the number is 27
the number is 49

Found 2 solutions by mananth, rfer:
Answer by mananth(16946)

the number is 49

Answer by rfer(16322)

3/7x=21
x=21*7/3
x=147/3
x=49